OpenStudy (abhisar):

two weights w1 and w2 are connected by a light thread which passes over a light smooth pulley if the pulley is raised upwards with an acceleration equal to g , then the tension in the thread will be @Kainui @nincompoop @iambatman @ParthKohli @skullpatrol

OpenStudy (abhisar):

Answer is \(\huge \sf \frac{4W_1W_2}{W_1+W_2}\) http://prntscr.com/4ew4mk

OpenStudy (midhun.madhu1987):

I was not aware of this...

OpenStudy (abhisar):

It's ok

OpenStudy (anonymous):

Drawing that goes with it?

OpenStudy (abhisar):

yes

OpenStudy (anonymous):

Oh it's not loading for me

OpenStudy (abhisar):

one min

OpenStudy (abhisar):

Can u explain this ?

OpenStudy (anonymous):

I'm not sure what the problem is here, it's just using newtons second law, pretty simple :P

OpenStudy (abhisar):

yeah :p

OpenStudy (anonymous):

Specific part? Or just not sure what's going on in general?

OpenStudy (abhisar):

Specific part

OpenStudy (abhisar):

\(\sf m_2g-T=m_2(g-a)\)

OpenStudy (abhisar):

Not able to understand the sign conventions !

OpenStudy (anonymous):

Look at the acceleration and gravity

OpenStudy (abhisar):

@No.name

OpenStudy (anonymous):

Wait reading the question!

OpenStudy (abhisar):

ok

OpenStudy (anonymous):

Is it quite urgent , because in 2 hours my exam is going to start , i am studying for it

OpenStudy (abhisar):

It's ok .....

OpenStudy (anonymous):

Yeah i am getting the same answer

OpenStudy (abhisar):

How ?

OpenStudy (abhisar):

@kropot72

OpenStudy (nincompoop):

do you still need help?

OpenStudy (nincompoop):

try breaking it down into parts

OpenStudy (abhisar):

@nincompoop yes i still need help...u there ?

OpenStudy (abhisar):

@compphysgeek @RaphaelFilgueiras

OpenStudy (abhisar):

I have this solution given, can u guys xplain it ?

OpenStudy (anonymous):

m2.g-T=m2(g-a), but g=a then T=m2.g i don't get this line

OpenStudy (anonymous):

the pseudo g, you can feel this effect when the elevator is going up dont tou feel your feet pressing the floor ?it's that effect

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @RaphaelFilgueiras m2.g-T=m2(g-a), but g=a then T=m2.g i don't get this line \(\color{blue}{\text{End of Quote}}\) Same here !

OpenStudy (abhisar):

I don't understand the sign convention they have used.

OpenStudy (anonymous):

is this the saraeva book?

OpenStudy (abhisar):

Idk...found it on internet...... I hv onemore solution from other source.......would u like to see it ?

OpenStudy (nincompoop):

what part of the solution you don't understand?

OpenStudy (abhisar):

How they deduced eq 1 and eq 2

OpenStudy (anonymous):

they actaully thought that m2 mass is greater than m1 so for m1 tension in the string is greater than the weight so they took T-m1 and for m2 ,its weight is greater than tension,so they took m2-T.. I THINK SOO BUT NOT SURE

OpenStudy (nincompoop):

are you familiar with the newton's laws?

OpenStudy (abhisar):

yes i am..

OpenStudy (nincompoop):

can you break the illustration into components according to what you know? forget the picture you were given, I want to know how you will make your own diagram.

OpenStudy (abhisar):

yes, ok let's suppose the lift is not moving and it is stationary....then it should be T-m1g = -m1a [if m1 > m2] T-m2g = m2a

OpenStudy (abhisar):

correct ?

OpenStudy (abhisar):

@nincompoop ?

OpenStudy (nincompoop):

when stationary, we use the first law T +(-w1) = 0 T = w1

OpenStudy (abhisar):

No i didn't said that the pulley is stationary.....i said the lift is stationary...!

OpenStudy (abhisar):

There are two simultaneous acceleration working according to the question...

OpenStudy (abhisar):

One is accn due to normal pulley thing and the other is due to the accelerated motion of lift inside which the pulley is kept.

OpenStudy (abhisar):

See this diagram, it represents the question

OpenStudy (abhisar):

The pulley is raised upwards with an acceleration=g

OpenStudy (nincompoop):

i see it now

OpenStudy (abhisar):

It's like raising a beam balance with two different weights on it

OpenStudy (nincompoop):

I understand that now. look at your equation when the pulley is not being raised

OpenStudy (abhisar):

No, w1 = m1g

OpenStudy (abhisar):

m1a is the accelaration by which the block will move downwards.

OpenStudy (abhisar):